# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v1

@@ 1 Introduction
Our objective is to accurately infer a posterior distribution over physically meaningful latent variables, to reconstruct the structural response, quantify the associated uncertainty, and predict the damage and degradation condition of the system in previously unseen conditions. This is achieved using a limited number of noisy measurements of the structural response, domain and class variables. Due to the influence of the domain, class, and other unknown confounding factors, this will generally be an ill-posed inverse problem that requires learning a disentangled representation [Locatello2019, Bengio2014] of the different generative factors. This task is further complicated by the limitations of physics-based models, which often represent structures under idealized nominal conditions and disregard the influence of environmental and operational variability, damage, and degradation. Most computational models of physical systems will contain simplifications and approximations due to lack of knowledge about certain aspects of the underlying physical process and to ensure computational tractability. Reducing this epistemic uncertainty is often infeasible due to cost or time constraints. As a result, only a partial description of the physical system is available in practical applications.
@@
  • Accurately quantify uncertainties by preventing the data-driven components of the decoder from compensating for all discrepancies between the physics-based model prediction and the measured response.
@@ 2 Background
This section aims to clarify the terminology and notation used throughout this text, summarize the necessary background, and illustrate the challenges that the proposed approach aims to address. In what follows, bold capital and lower case symbols denote matrix and vector quantities respectively. Light symbols denote scalars. Latent variables that are not directly observed and must be inferred from data are denoted as $z$, while $\phi$, $\theta$ and $\psi$ denote encoder, decoder and auxiliary regressor/classifier parameters, respectively. The symbols $x$, $c$ and $y$ denote the response, domain, and class observables, jointly referred to as the modalities of a given physical system. When used as a subscript these symbols denote quantities that belong to a particular modality. As an example, $\bm{z}_{\mathrm{y}}$ denotes a set of latent variables that encode information about the class of a physical system. A distinction is made between the underlying generative factors $s$ that determine the characteristics of the observed data, and the latent variables $z$, i.e. the learned representation of the generative factors.
@@ 2.1 Problem setting
To highlight the intended application setting, three examples are presented in Figure 1 consisting of a Euler-Bernoulli beam, a damped oscillator, and population of bridges. In each example the available physics-based model fails to account for domain and class influences that are present in the measured response signals. In all three examples, the domain and class observables provide valuable information about the system that is necessary for accurately inferring a distribution over parameters of the physics-based model and reconstructing the response of the physical system. Furthermore, the class observables are significantly harder to measure and will not be available for future experiments. The damped oscillator and bridge examples also include unknown confounding influences, for which neither observations nor a physical description are available. The three examples are summarized below:
@@
Figure 1: Illustrative examples of the problem setting: a) Euler-Bernoulli beam, b) Damped oscillator, and c) One member of a population of bridges. The objective is to learn components of the measured response (bottom row) that are not explicitly included in the nominal physics-based model (top row) using observations of related quantities.
@@
    Euler-Bernoulli beam: In this example, the available physics-based model of a beam assumes simply supported boundary conditions and a point load acting on an unknown position. Noisy measurements of the displacement field are obtained from a set of sensors, equally spaced along the length of the beam. The domain influence is introduced as a dependence of the rotational stiffness of the right support on the temperature, and the class influence is taken as damage causing a reduction in the vertical stiffness of the right support that varies between experiments.
@@
    Damped oscillator: Noisy displacement time-series are obtained from multiple experiments, where a mass-spring-dashpot system is deflected from the equilibrium position and released to perform a harmonic oscillation. The impact of damping on the motion is neglected in the physics-based model and must be inferred from additional observations of properties of the medium. The spring stiffness is taken to depend on the ambient temperature, and there is unknown variability in the initial displaced position. The damping, temperature and variability in initial position are taken as class, domain and unknown confounding influences, respectively.
@@ 2.3 The Variational Autoencoder
To remedy the lack of domain knowledge, data-driven models based on machine learning techniques have emerged as an alternative to physics-based models, where the unknown physical process is learned from measurements using flexible parametrized approximations. VAE [Kingma2022, Kingma2019] are a popular data-driven approach for learning a joint distribution of data and the latent variables that are assumed to have generated the data using amortized Variational Inference (VI) [Blei2017]. In VAE, the per-datapoint posterior distribution is approximated using a parameterized family of distributions, where the optimal parameters are obtained by minimizing the Kullback-Leibler divergence (KLD) between the true and approximate posteriors. The VAE is composed of an encoder network $q_{\mathrm{\bm{\phi}}}(\bm{z}|\bm{x})$ and a decoder network $p_{\theta}(\bm{x}|\bm{z})$, parameterized by $\phi$ and $\theta$ respectively, where $\bm{z}$ denotes latent variables that can not be observed directly and must be inferred from measurements. The encoder is typically implemented as a feed-forward NN that maps the inputs $\bm{x}$ to a conditional density over latent variables $\bm{z}$. The decoder network $p_{\mathrm{\bm{\theta}}}(\bm{x}|\bm{z})$ works in the opposite direction by approximating the density of $\bm{x}$ conditioned on $\bm{z}$. The training process for VAE consists of simultaneously optimizing the parameters of the decoder that reconstructs the observations samples of the latent variables, and the encoder that maps inputs to a posterior distribution over these latent variables. Optimization is performed by maximizing a lower bound on the marginal likelihood of the data known as the Evidence Lower BOund (ELBO), denoted as $\mathcal{L}_{\mathrm{VAE}}$ in Equation 2.1, where $\mathbb{E}_{p(\cdot)}\left[f(\cdot)\right]$ is the expectation of a function $f(\cdot)$ over a distribution $p(\cdot)$. Sampling $z\sim q_{\phi}(\bm{z}|\bm{x})$ and evaluating the decoder yields samples from the learned distribution of the data, which in the context of civil and mechanical structural systems can be used for downstream tasks such as remaining useful life assessment.
@@ 2.4 Challenges in combining physics-based and data-driven components in VAE
A straightforward approach to account for epistemic uncertainty in a data-driven manner would be to approximate the measured response $\bm{x}$ as the sum of the physics-based model [^1] It is assumed that the gradients of the function with respect to the inputs can be evaluated efficiently to obtain a computationally tractable optimization problem. $f(\bm{z}_{\mathrm{x}})$ and a trainable NN-based function $g_{\bm{\theta}}(\cdot)$, where the latter corrects the discrepancies between the physics-based model predictions and measurements. This type of hybrid model is referred to as a residual model. It is obvious that parametrizing the data driven component of the residual model as $g_{\bm{\theta}}(\bm{z}_{\mathrm{x}})$ is not feasible when it is required that $\bm{z}_{\mathrm{x}}$ is interpretable: The resulting hybrid generative model has a posterior distribution $p_{\mathrm{\theta}}(\bm{z}_{\mathrm{x}}|\bm{x})$, where the latent variables $\bm{z}_{\mathrm{x}}$ are the input to a coupled physics-based and data-driven model, and thus no longer physically meaningful. Instead, the latent space can be partitioned as $(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\in\bm{z}$ and the data driven component parametrized as $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, where $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ are physically meaningless latent variables intended to capture variability in the measured response due to the influence of the domain and the class. Assuming that the remaining aleatory uncertainties (e.g. caused by measurement noise) are independent of the signal being measured and can be sufficiently modeled as independent and identically distributed (i.i.d.) samples of Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}$, the response measurements can be expressed as:
@@
The observables $\bm{c}$ and $\bm{y}$ will be used to ensure that the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ encode information about the domain and class of the structure by simultaneously the training auxiliary tasks $r_{\mathrm{c}}(\bm{c}|\bm{z}_{\mathrm{c}})$ and $r_{\mathrm{y}}(\bm{y}|\bm{z}_{\mathrm{y}})$. The resulting hybrid generative model $p_{\bm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ can be coupled with variational posteriors $q_{\mathrm{\phi}_{\mathrm{x}}}(\bm{z}_{\mathrm{x}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{x})$ to yield an architecture similar to VAE. This is the architecture derived in Section 3, without the additional constraints and regularization. This parametrization circumvents the previous issue, but does not ensure that the known physics are utilized by the model, or that $\bm{z}_{\mathrm{x}}$ is physically meaningful. Without further regularization, the model can learn combinations of arbitrary predictions from the physics-based and data-driven components $f(\bm{z}_{\mathrm{x}})$ and $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ that sum to an accurate prediction. To see why, it is sufficient to consider the form of the objective given in Equation 2.1. Both the encoder and decoder are aligned in the task of maximizing the reconstruction term $\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}|\bm{z})\right]$, and the data-driven components of the model will account for any discrepancy between the physics-based model prediction and measurements up to some level of noise. This results in an entangled representation, where the data-driven components override the known physics and the physics-grounded latent variables loose their physical meaning.
This issue is illustrated in Figure 2 using the Euler-Bernoulli beam case study shown in Figure 1(a). Further details of the case study are provided in Section 5.1. In this example, a VAE trained on a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{c}_{i},\bm{y}_{i})\}_{i=1}^{N}$ is evaluated on a new set of input measurements $\bm{x}$, generated from the ground truth data generating process by linearly varying the Young’s modulus $E$. It can be seen that the data-driven component of the decoder overrides the known physics. Furthermore, the physics-based component $\hat{\bm{x}}_{\mathrm{p}}$ of the predicted response $\hat{\bm{x}}$ is the result of variation of both the Young’s modulus $E$ and the position of the load $x_{\mathrm{F}}$, despite the fact that the latter should remain constant. Conversely, the data-driven component $\hat{\bm{x}}_{\mathrm{d}}$ counteracts the variation of $x_{\mathrm{F}}$, resulting in an accurate combined prediction. This effect can be inconsistent and will depend on the details of the neural network components in the model and the physics of the problem at hand. Here it is replicated consistently by reducing the learning rate of the encoder $q_{\mathrm{\phi}_{\mathrm{x}}}(\bm{z}_{\mathrm{x}}|\bm{x})$ by a factor $10$.
@@
Figure 2: Demonstration of the data-driven component of the decoder $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ overriding the physics-based model $f(\bm{z}_{\mathrm{x}})$. The effect of varying the Young’s modulus $E$ should be described by the known physics, but is instead captured by both the physics-based and data-driven components.
The results presented in Figure 2 are obtained under the assumption of a factorized variational posterior. Ideally, a single encoder with shared parameters $q_{\mathrm{\phi}}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})$ would be used for the three subsets of the latent space $\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. However, using a shared encoder leads to further degradation of the performance as noted by [Ilse2020]. This example demonstrates how the interaction between the physics-based and data-driven components, which determines the resulting learned representation, depends on the capacity and flexibility of the individual machine learning components. Standard VAE offer no mechanism to regulate this interaction, and are therefore unable to guarantee that the physics will be utilized correctly in the presence of flexible NN-based decoder components.
@@ 3.1 Detailed description of the model
It is assumed that three generative factors, the underlying physics of the structure, the domain, and the class, contribute to the measured response $\bm{x}$. Conversely, the latent variables are partitioned into subsets [^2] It is emphasized that the separation of the latent variables is only semantic and used for clarity. In practice, they can be the output of a single encoder with shared parameters. $(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\in\bm{z}$. The latent variables are the input to the hybrid probabilistic decoder $p_{\theta}(\bm{x}|\bm{z})$, wherein a NN-based function $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ accounts for discrepancies between the measured response $\bm{x}$ and physics-based model prediction $f(\bm{z}_{\mathrm{x}})$. To ensure that information relevant to the domain and class is encoded in the corresponding subsets $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, we utilize two auxiliary decoders $r_{\psi_{\mathrm{c}}}(\bm{c}|\bm{z}_{c})$ and $r_{\psi_{\mathrm{y}}}(\bm{y}|\bm{z}_{y})$. The latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ are assigned conditional prior distributions $p_{\theta_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{c})$ and $p_{\theta_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{y})$ respectively, while the physics-grounded latent variables $\bm{z}_{\mathrm{x}}$ will be assigned a distribution $p(\bm{z}_{\mathrm{x}})$ based on the available prior knowledge. A simplified schematic illustration of the architecture is provided in Figure 3(a).
@@
The quantities described in Equation 3.1 are optimized indirectly through a latent bottleneck structure [Tishby2000, Alemi2019, Fischer2020, Moyer2018] combined with adversarial training, as described in the following informal sketch. Note that in the inference model shown in Figure 3(b), the latent variables $\bm{z}_{\mathrm{y}}$ do not depend directly on $\bm{y}$ (and analogously for $\bm{c}$). This is the conditional independence assumption typically used in the Information Bottleneck framework. Intuitively, the encoder is forced to distill the relevant information in $\bm{x}$ that is necessary for reconstructing $\bm{c}$ and $\bm{y}$ into the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. This results in the maximization of the $I_{\mathrm{\phi}}(\bm{c};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{y};\bm{z}_{\mathrm{y}})$ terms in Equation 3.1 during training. The additional requirement of minimizing $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{y}})$ can be satisfied by introducing a Gradient Reversal Layer (GRL) [Ganin2015] at the input of the data-driven decoder component $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$. During optimization, the gradient signal propagated backwards from $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ to the encoder is scaled by $-\lambda$, while the forward pass remains unchanged. Therefore, the GRL can be thought of as a pseudo function $R_{\mathrm{\lambda}}(\bm{z})$ such that $R_{\mathrm{\lambda}}(\bm{z})=\bm{z}$ and $\frac{\mathrm{d}R_{\mathrm{\lambda}}}{\mathrm{d}\bm{z}}=-\lambda\bm{I}$. Positive values of $\lambda$ correspond to adversarial training. Conversely, negative values make the training collaborative. The absolute value of $\lambda$ determines the strength of the adversarial or collaborative objective, with larger values corresponding to a stronger regularization effect. By turning the decoder $p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ into an adversary, the GRL penalizes information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ proportionally to its contribution to the reconstruction of $\bm{x}$, biasing the encoder towards representations that are minimally informative about $\bm{x}$.
@@
Rearranging the terms in Equation 3.2, noting that the generative model factorizes as $p(\bm{x},\bm{c},\bm{y},\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})=p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})p(\bm{z}_{\mathrm{x}})p_{\mathrm{\theta}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{x})p_{\mathrm{\theta}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{y})p(\bm{c})p(\bm{y})$, yields the following expression for the lower bound:

### v2

@@ 1 Introduction
Our objective is to accurately infer a posterior distribution over physically meaningful latent variables, to reconstruct the structural response, quantify the associated uncertainty, and predict the damage and degradation condition of the system in previously unseen conditions. This is achieved using a limited number of noisy measurements of the structural response, domain and class variables. Due to the influence of the domain, class, and other unknown confounding factors, this will generally be an ill-posed inverse problem that requires learning a disentangled representation [Bengio2014] of the different generative factors. This task is further complicated by the limitations of physics-based models, which often represent structures under idealized nominal conditions and disregard the influence of environmental and operational variability, damage, and degradation. Most computational models of physical systems will contain simplifications and approximations due to lack of knowledge about certain aspects of the underlying physical process and to ensure computational tractability. Reducing this epistemic uncertainty is often infeasible due to cost or time constraints. As a result, only a partial description of the physical system is available in practical applications.
@@
  • Improve uncertainty quantification by preventing the data-driven components of the decoder from compensating for all discrepancies between the physics-based model prediction and the measured response.
@@ 2 Background
This section aims to clarify the terminology and notation used throughout this text, summarize the necessary background, and illustrate the challenges that the proposed approach aims to address. In what follows, bold capital and lower case symbols denote matrix and vector quantities respectively. Light symbols denote scalars. Latent variables that are not directly observed and must be inferred from data are denoted as $z$, while $\phi$, $\theta$ and $\psi$ denote encoder, decoder and auxiliary regressor/classifier parameters, respectively. The symbols $x$, $c$ and $y$ denote the response, domain, and class observables, jointly referred to as the modalities of a given physical system. When used as a subscript these symbols denote quantities that belong to a particular modality. As an example, $\bm{z}_{\mathrm{y}}$ denotes a set of latent variables that encode information about the class of a physical system. Throughout this text, $\mathcal{N}$ denotes the univariate or multivariate normal distribution parametrized by the mean and a scalar variance or matrix covariance respectively, and $\mathcal{U}$ denotes the uniform distribution parametrized by the lower and upper bound. The expectation of a function $f(\cdot)$ over a distribution $p(\cdot)$ is denoted as $\mathbb{E}_{p(\cdot)}\left[f(\cdot)\right]$. Finally, a distinction is made between the underlying generative factors $s$ that determine the characteristics of the observed data, and the latent variables $z$, i.e. the learned representation of the generative factors.
@@ 2.1 Problem setting
To highlight the intended application setting, three examples are presented in Figure 1 consisting of a beam, an oscillator, and population of bridges. In each example the available physics-based model fails to account for domain and class influences that are present in the measured response signals. In all three examples, the domain and class observables provide valuable information about the system that is necessary for accurately inferring a distribution over parameters of the physics-based model and reconstructing the response of the physical system. Furthermore, the class observables are significantly harder to measure and will not be available for future experiments. The oscillator and bridge examples also include unknown confounding influences, for which neither observations nor a physical description are available. The three examples are summarized below:
@@
Figure 1: Illustrative examples of the problem setting: a) Beam, b) Oscillator, and c) One member of a population of bridges. The objective is to learn components of the measured response (bottom row) that are not explicitly included in the nominal physics-based model (top row) using observations of related quantities.
@@
    Beam: In this example, the available physics-based model of a beam assumes simply supported boundary conditions and a point load acting on an unknown position. Noisy measurements of the displacement field are obtained from a set of sensors, equally spaced along the length of the beam. The domain influence is introduced as a dependence of the rotational stiffness of the right support on the temperature, and the class influence is taken as damage causing a reduction in the vertical stiffness of the right support that varies between experiments.
@@
    Oscillator: Noisy displacement time-series are obtained from multiple experiments, where a mass-spring-dashpot system is deflected from the equilibrium position and released to perform a harmonic oscillation. The impact of damping on the motion is neglected in the physics-based model and must be inferred from additional observations of properties of the medium. The spring stiffness is taken to depend on the ambient temperature, and there is unknown variability in the initial displaced position. The damping, temperature and variability in initial position are taken as class, domain and unknown confounding influences, respectively.
@@ 2.3 The Variational Autoencoder
To remedy the lack of domain knowledge, data-driven models based on machine learning techniques have emerged as an alternative to physics-based models, where the unknown physical process is learned from measurements using flexible parametrized approximations. VAE [Kingma2022, Kingma2019] are a popular data-driven approach for learning a joint distribution of data and the latent variables that are assumed to have generated the data using amortized Variational Inference (VI) [Blei2017]. In VAE, the per-datapoint posterior distribution is approximated using a parametrized family of distributions, where the optimal parameters are obtained by minimizing the Kullback-Leibler divergence (KLD) between the true and approximate posteriors. The VAE is composed of an encoder network $q_{\mathrm{\bm{\phi}}}(\bm{z}|\bm{x})$ and a decoder network $p_{\theta}(\bm{x}|\bm{z})$, parametrized by $\phi$ and $\theta$ respectively, where $\bm{z}$ denotes latent variables that can not be observed directly and must be inferred from measurements. The encoder is typically implemented as a feed-forward NN that maps the inputs $\bm{x}$ to a conditional density over latent variables $\bm{z}$. The decoder network $p_{\mathrm{\bm{\theta}}}(\bm{x}|\bm{z})$ works in the opposite direction by approximating the density of $\bm{x}$ conditioned on $\bm{z}$. The training process for VAE consists of simultaneously optimizing the parameters of the decoder that reconstructs the observations samples of the latent variables, and the encoder that maps inputs to a posterior distribution over these latent variables. Optimization is performed by maximizing a lower bound on the marginal likelihood of the data known as the Evidence Lower BOund (ELBO), denoted as $\mathcal{L}_{\mathrm{VAE}}$ in Equation 2.1. Sampling $z\sim q_{\phi}(\bm{z}|\bm{x})$ and evaluating the decoder yields samples from the learned distribution of the data, which in the context of civil and mechanical structural systems can be used for downstream tasks such as remaining useful life assessment.
@@ 2.4 Challenges in combining physics-based and data-driven components in VAE
A straightforward approach to account for epistemic uncertainty in a data-driven manner would be to approximate the measured response $\bm{x}$ as the sum of the physics-based model [^1] It is assumed that the gradients of the function with respect to the inputs can be evaluated efficiently to obtain a computationally tractable optimization problem. $f(\bm{z}_{\mathrm{x}})$ and a trainable NN-based function $g_{\bm{\theta}}(\cdot)$, where the latter corrects the discrepancies between the physics-based model predictions and measurements. This type of hybrid model is referred to as a residual model. However, parametrizing the data driven component of the residual model as $g_{\bm{\theta}}(\bm{z}_{\mathrm{x}})$ is not feasible when it is required that $\bm{z}_{\mathrm{x}}$ is interpretable: The resulting hybrid generative model has a posterior distribution $p_{\mathrm{\theta}}(\bm{z}_{\mathrm{x}}|\bm{x})$, where the latent variables $\bm{z}_{\mathrm{x}}$ are the input to a coupled physics-based and data-driven model, and thus no longer physically meaningful. Instead, the latent space can be partitioned as $(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\in\bm{z}$ and the data driven component parametrized as $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, where $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ are physically meaningless latent variables intended to capture variability in the measured response due to the influence of the domain and the class. Assuming that the remaining aleatory uncertainties (e.g. caused by measurement noise) are independent of the signal being measured and can be sufficiently modeled as independent and identically distributed (i.i.d.) samples of Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}$, the response measurements can be expressed as:
@@
The observables $\bm{c}$ and $\bm{y}$ can be used to ensure that the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ encode information about the domain and class of the structure by simultaneously training auxiliary tasks $r_{\mathrm{c}}(\bm{c}|\bm{z}_{\mathrm{c}})$ and $r_{\mathrm{y}}(\bm{y}|\bm{z}_{\mathrm{y}})$. The resulting hybrid generative model $p_{\bm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ can be coupled with variational posteriors $q_{\mathrm{\phi}_{\mathrm{x}}}(\bm{z}_{\mathrm{x}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{x})q_{\mathrm{\phi}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{x})$ to yield an architecture similar to VAE. This is the architecture derived in Section 3, without the additional constraints and regularization. This parametrization circumvents the previous issue, but does not ensure that the known physics are utilized by the model, or that $\bm{z}_{\mathrm{x}}$ is physically meaningful. Without further regularization, the model can learn combinations of arbitrary predictions from the physics-based and data-driven components $f(\bm{z}_{\mathrm{x}})$ and $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ that sum to an accurate prediction. To see why, it is sufficient to consider the form of the objective given in Equation 2.1. Both the encoder and decoder are aligned in the task of maximizing the reconstruction term $\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}|\bm{z})\right]$, and the data-driven components of the model will account for discrepancies between the physics-based model prediction and measurements up to some level of noise. This results in an entangled representation, where the data-driven components override the known physics and the physics-grounded latent variables loose their physical meaning. This issue is illustrated in Figure 2 using the beam case study shown in Figure 1(a). Further details of the case study are provided in Section 5.1. In this example, a VAE trained on a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{c}_{i},\bm{y}_{i})\}_{i=1}^{N}$ is evaluated on a new set of input measurements $\bm{x}$, generated from the ground truth data generating process by linearly varying the position of the load $x_{\mathrm{F}}$. It can be seen that the effect of this variation on the measured response is largely captured by the data-driven component of the decoder, which overrides the known physics, despite the fact that the physics-based model includes the load position as an input parameter. The extent to which the data-driven components override the known physics can be inconsistent and hard to predict, and will depend on the neural network architectures and the physics of the problem at hand.
@@
Figure 2: Demonstration of the data-driven component of the decoder $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ overriding the physics-based model $f(\bm{z}_{\mathrm{x}})$. The effect of varying the position of the load $x_{\mathrm{F}}$ should be described by the known physics, but is instead captured by the data-driven components.
The results presented in Figure 2 are obtained under the assumption of a factorized variational posterior. Ideally, a single encoder with shared parameters $q_{\mathrm{\phi}}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})$ would be used for the three subsets of the latent space $\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. However, using a shared encoder leads to further degradation of the performance as noted by [Ilse2020]. This example demonstrates how the interaction between the physics-based and data-driven components, which determines the resulting learned representation, depends on the capacity and flexibility of the individual machine learning components. Standard VAE offer no mechanism to control this interaction, and are therefore unable to guarantee that the physics will be utilized correctly in the presence of flexible NN-based decoder components.
@@ 3.1 Detailed description of the model
It is assumed that three generative factors, the underlying physics of the structure, the domain, and the class, contribute to the measured response $\bm{x}$. Conversely, the latent variables are partitioned into subsets [^2] It is emphasized that the separation of the latent variables is only semantic and used for clarity. In practice, they can be the output of a single encoder with shared parameters. $(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\in\bm{z}$. The latent variables are the input to the hybrid probabilistic decoder $p_{\theta}(\bm{x}|\bm{z})$, wherein a NN-based function $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ accounts for discrepancies between the measured response $\bm{x}$ and physics-based model prediction $f(\bm{z}_{\mathrm{x}})$. To ensure that information relevant to the domain and class is encoded in the corresponding subsets $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, we utilize two auxiliary decoders $r_{\psi_{\mathrm{c}}}(\bm{c}|\bm{z}_{c})$ and $r_{\psi_{\mathrm{y}}}(\bm{y}|\bm{z}_{y})$. The latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ are assigned conditional prior distributions $p_{\theta_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{c})$ and $p_{\theta_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{y})$ respectively, while the physics-grounded latent variables $\bm{z}_{\mathrm{x}}$ are assigned a distribution $p(\bm{z}_{\mathrm{x}})$ based on the available prior knowledge. A schematic illustration of the architecture is provided in Figure 3(a).
@@
The quantities described in Equation 3.1 are optimized indirectly through a latent bottleneck structure [Tishby2000, Alemi2019, Fischer2020, Moyer2018] combined with adversarial training, as described in the following informal sketch. Note that in the inference model shown in Figure 3(b), the latent variables $\bm{z}_{\mathrm{y}}$ do not depend directly on $\bm{y}$ (and analogously for $\bm{c}$). This is the conditional independence assumption typically used in the Information Bottleneck framework. Intuitively, the encoder is forced to distill the relevant information in $\bm{x}$ that is necessary for reconstructing $\bm{c}$ and $\bm{y}$ into the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. This results in the maximization of the $I_{\mathrm{\phi}}(\bm{c};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{y};\bm{z}_{\mathrm{y}})$ terms in Equation 3.1 during training. The additional requirement of minimizing $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{y}})$ can be satisfied by introducing a Gradient Reversal Layer (GRL) [Ganin2015] at the input of the data-driven decoder component $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$. During optimization, the gradient signal propagated backwards from $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ to the encoder is scaled by $-\lambda$, while the forward pass remains unchanged. Therefore, the GRL can be thought of as a pseudo function $R_{\mathrm{\lambda}}(\bm{z})$ such that $R_{\mathrm{\lambda}}(\bm{z})=\bm{z}$ and $\frac{\mathrm{d}R_{\mathrm{\lambda}}}{\mathrm{d}\bm{z}}=-\lambda\bm{I}$. Positive values of $\lambda$ correspond to adversarial training. Conversely, negative values make the training “collaborative”. The absolute value of $\lambda$ determines the strength of the adversarial or collaborative objective, with larger values corresponding to a stronger regularization effect. By turning the decoder $p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ into an adversary, the GRL penalizes information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that contributes to the reconstruction of $\bm{x}$, biasing the encoder towards representations that are minimally informative about $\bm{x}$.
@@
Rearranging the terms in Equation 3.2, noting that the generative model factorizes as $p(\bm{x},\bm{c},\bm{y},\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})=p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})p(\bm{z}_{\mathrm{x}})p_{\mathrm{\theta}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{c})p_{\mathrm{\theta}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{y})p(\bm{c})p(\bm{y})$, yields the following expression for the lower bound: